OpenStudy (anonymous):

simplify the expression. (3-i)-(2+6i) A.1-7i B.-1+7i C.-6i D.5+5i

OpenStudy (solomonzelman):

factor it out. Foil!

OpenStudy (anonymous):

my guess is A

OpenStudy (shamil98):

This isn't factoring, solomon.. 3 - i - 2 - 6i 1 - 7i

OpenStudy (solomonzelman):

Foiling

OpenStudy (shamil98):

you distribute the negative one to (2 + 6i) then you just combine like terms, solomon

OpenStudy (solomonzelman):

Oh, didn't see the minus, MY BAD

OpenStudy (anonymous):

\[(3-i)-(2+6i) = (3-i)+(-2-6i)\] Place common symbols together \[3-2-i-6i=1-7i\]

OpenStudy (anonymous):

:D i was right thanks